OpenStudy (gentorr):

Light will bend as it slows down in space. speeds up in space. changes speeds passes from one medium to another. there are 2 answers i think i know one of them which is d

OpenStudy (gentorr):

@inkyvoyd

OpenStudy (gentorr):

@linn99123

OpenStudy (gentorr):

@TheEdwardsFamily

OpenStudy (gentorr):

@Demonx341

OpenStudy (gentorr):

sorry just asking for more people to come

OpenStudy (theedwardsfamily):

It's fine. I'm just a little confused with this question :P

OpenStudy (gentorr):

me too

OpenStudy (gentorr):

i only ask a question if im really confused

OpenStudy (theedwardsfamily):

We have two kinds of light speed: the speed of photons C whose flight path is unencumbered and the average speed of photons c < C whose flight path is encumbered by particles. The average speed through media filled with particles is slower than the speed without particles in the way because of dwell time, d. As photons couple with electrons of the media, they are destroyed and their energy raises the electron energies. New photons emerge when the electrons pop back to a lower energy level. All that processing, that coupling and uncoupling, takes time, dwell time. So if space S is crossed in t = S/C time without particles to couple with, the total time to cross S with dwell time will be T = SUM(d) + t > t. It simply takes more time on average to cross the same space S when the space is filled with stuff to run into. And there we are: c = S/T < S/t = C so that the average speed through a medium will always...always...be slower than through a vacuum, space without particles to run into. And not all media are created equal. Some will have more particles, some will have less, some will be more inclined to couple, some will be less. So we find that some material, like air, has average speed c that's almost the same as C. And of course, lots of things are totally opaque; so the speed is c = 0. NOTE: We know that light bends when going through lenses and such; but we really don't know why. In HS we were taught a myth that wave fronts had to stay together and that's why the rays bent, to keep the wave fronts together as they go from medium to medium. But that turns out to be a myth; there is absolutely no physical reason why wavefronts must stay intact. In fact the concept of rays of light is faulty. To date the best we can come up with that sort of makes physical sense is the sum of all probable paths. When we do the math, really heavy math, the vectors do form what we can define as rays of light.

OpenStudy (theedwardsfamily):

Got that from yahoo answers. It might help you

OpenStudy (gentorr):

yea i saw just didnt read it but i'll read it know

OpenStudy (theedwardsfamily):

Ok. :D

OpenStudy (gentorr):

thank you

OpenStudy (theedwardsfamily):

D should be right tho

OpenStudy (theedwardsfamily):

You're welcome! ^_^

OpenStudy (gentorr):

i think it might be C. it changes speed because of this part We have two kinds of light speed: the speed of photons C whose flight path is unencumbered and the average speed of photons c < C whose flight path is encumbered by particles.

OpenStudy (gentorr):

what do you think

OpenStudy (theedwardsfamily):

hm...yeah I suggest c too. Makes sense.

OpenStudy (gentorr):

okay thank

OpenStudy (theedwardsfamily):

you welcome! Medal 4 you 2 :P

OpenStudy (gentorr):

why and thank you -_O

OpenStudy (theedwardsfamily):

Lol because you were figuring it out too

OpenStudy (gentorr):

okay well thank you

OpenStudy (theedwardsfamily):

np c;

OpenStudy (gentorr):

if i have another question can you help

OpenStudy (theedwardsfamily):

Sure! I'll try

OpenStudy (gentorr):

kk

OpenStudy (theedwardsfamily):

Just tag me if you need me :)

OpenStudy (gentorr):

kk